\begin{thebibliography}{9}

\bibitem{AZ}T.  Amdeberhan and D. Zeilberger,
$q$-Apery irrationality proofs by $q$-WZ pairs, Adv. Appl. Math.
 20 (1998) 275--283.


     \bibitem{And2}G.E. Andrews, Ramanujan's ``lost" book. I.
     Partial $\theta$-functions, Adv. Math. 41 (1981) 137--172.

\bibitem{And3} G.E. Andrews,
     Pfaff's Method (I): the Mills-Robbins-Rumsey determinant.
      Discrete Math. 193 (1998) 43--60.


\bibitem{And4} G.E. Andrews,
Pfaff's Method (II): diverse applications. J. Computational and
Appl. Math. 68 (1996) 15--23.

\bibitem{And6} G.E. Andrews,
Pfaff's Method (III): comparison with the WZ method. Electronic J.
Combin. (Foata Festschrift)    3(2) (1996)   R21.

   \bibitem{GS06} G.E. Andrews and S. Ole Warnaar, The product of
   partial theta functions. Adv. Applied Math.
   39 (2007) 116--120.

   \bibitem{Bailey} W.N. Bailey, Identities of the Rogers-Ramanujan
   type, Proc. London Math. Soc. 50 (1949) 1--10.


\bibitem{CHM2}
 W.Y.C. Chen, Q.H. Hou and Y.P. Mu,  Nonterminating
 basic hypergeometric series
  and the q-Zeilberger algorithm,
   Proc. Edinburgh Math. Soc.  to appear.


   \bibitem{GR2004} G. Gasper and M. Rahman,
   Basic Hypergeometric Series, Encyclopedia of Mathematics
   and its Applications, Vol. 35, second edition,
   Cambridge Univ. Press, Cambridge, MA, 2004.


   \bibitem{Gosper}
     R.W. Gosper, Decision procedure for indefinite
     hypergeometric summation, Proc. Natl. Acad. Sci. USA
       75(1) (1978) 40--42.

      \bibitem{Gessel}
     I.M. Gessel, Finding identities with the WZ mothed, J. Symbolic
     Computation  20 (1995) 537--566.


  \bibitem{Koepf03}
W. Koepf, Computer algebra algorithms for orthogonal polynomials and
special functions, In: Orthogonal Polynomials and Special Functions:
Leuven 2002, E. Koelink and W. Van Assche Eds., Lecture Notes Math.
 1817, 2003, pp. 1--24.

     \bibitem{Koor0}
     T.H. Koornwinder, On Zeilberger's algorithm and its $q$-analogue, J.
     Comput. Appl. Math. 48 (1993) 91--111.


   \bibitem{Koor}
   T.H. Koornwinder, Identities of nonterminating series by
  Zeilberger's algorithm, J. Comput. Appl. Math. {\bf99} (1998)
  449--461.

  \bibitem{Koorweb}T.H.  Koornwinder,
 {\tt http://staff.science.uva.nl/\~\,thk/specfun/compalg.html.}



  \bibitem{MZ}M. Mohammed and D. Zeilberger,
  The Markov-WZ method,  Electronic J. Combin.  11 (2004)
   R53.


    \bibitem{Paul} P. Paule, On identities of the Rogers-Ramanujan
    type, J. Math. Anal. Appl. 107 (1985) 255--284.

    \bibitem{PWZ}
    M. Petkov\v{s}ek, H.S. Wilf, and D. Zeilberger, A=B,
    A. K. Peters, Wellesley,
    MA, 1996.

    \bibitem{Ramanujan}
     S. Ramanujan, The ``Lost Notebook" and Other Unpublished Papers,
    Narosa Publishing House,
       New Delhi, 1988.



     \bibitem{SO2003} S.O. Warnaar, Partial theta functions. I.
   Beyond the lost book, Proc. London Math. Soc.  87
   (2003) 363--395.

   \bibitem{Vidunas02} R. Vidunas,
   A generalization of Kummer's identity,
 Rocky Mountain
J. Math.  32 (2002) 919--936.


\bibitem{WZ1}
H. S. Wilf and D. Zeilberger, Towards computaterized proofs of
identities, Bull. Amer. Math. Soc. 23(1) (1990) 77--83.

\bibitem{WZ2}
H.S. Wilf and D. Zeilberger, Rational functions certify
combinatorial identities, J. Amer. Math. Soc. 3(1) (1990) 147--158.


\bibitem{WZ3}
H.S. Wilf and D. Zeilberger, An algorithmic proof theory for
hypergeometric (ordinary and ``$q$") multisum/integral identities,
Invent. Math. 108 (1992) 575--633.




\bibitem{Zeilberger1} D. Zeilberger,
A fast algorithm for proving terminating hypergeometric identities,
Discrete Math. 80 (1990) 207--211.

\bibitem{Zeilberger2} D. Zeilberger,
Towards a WZ proof of Mehta's integral, SIAM J. Math. Anal. 25
 (1994)  812--814.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amssymb,amsfonts,amsmath,amsthm,cite}
\usepackage{epsfig}
\newtheorem{thm}{Theorem}[section]

\makeatletter \@addtoreset{equation}{section}
\def\qed{\hfill \rule{4pt}{7pt}}

\begin{document}

\begin{center}
{\bf\Large{The $q$-WZ Method for Infinite Series} }
\end{center}

\begin{center}
William Y. C. Chen$^1$ and Ernest X. W. Xia$^2$

Center for Combinatorics, LPMC-TJKLC\\
Nankai University\\
 Tianjin 300071, P. R. China

Email: $^1$chen@nankai.edu.cn,  $^2$xxwrml@mail.nankai.edu.cn

\end{center}

%===========================================================================
\noindent {\bf Abstract.} Motivated by the telescoping proofs of two
identities of Andrews and Warnaar, we find that  infinite
$q$-shifted factorials can be incorporated into the implementation
of the $q$-Zeilberger algorithm in the approach of  Chen, Hou and Mu
to prove nonterminating basic hypergeometric series identities. This
observation enables us to extend the $q$-WZ method to identities on
infinite series.
 As examples, we
will give the $q$-WZ pairs for some classical  identities such as
the $q$-Gauss sum, the $_6\phi_5$ sum,  Ramanujan's  $_1\psi_1$ sum
and Bailey's $_6\psi_6$ sum.

\noindent {\bf Keywords:} basic hypergeometric series, the
$q$-Gosper algorithm, the $q$-Zeilberger algorithm, the $q$-WZ
method.

\noindent {\bf AMS Classification:} 33D15, 33F10.
%===========================================================================

\section{Introduction}

Motivated by the telescoping proofs of two identities of Andrews and
Warnaar  \cite{GS06}, we find that infinite $q$-shifted factorials
can be incorporated into the implementation of the $q$-Zeilberger
algorithm as in the approach of Chen, Hou and Mu \cite{CHM2} to
prove nonterminating basic hypergeometric series identities. This
idea also leads to an extension of the $q$-WZ method to
nonterminating basic hypergeometric series identities.

 We will follow the
standard notation on $q$-series {\cite{GR2004}} and always assume
$|q|<1$. The $q$-shifted factorials $(a; q)_n$ and $(a;q)_\infty$
are defined as
\begin{eqnarray*}
(a;q)_n & = &\left\{
     \begin{array}{ll}
         1, \qquad & \mbox{if $n=0$}, \\ [8pt]
        (1-a) (1-aq) \cdots (1-aq^{n-1}), \qquad
              & \mbox{if $n\geq 1$},
     \end{array} \right. \\ [10pt]
     (a;q)_{-n}&=&\frac{1}{(aq^{-n};q)_n}, \qquad \mbox{if $n\geq
     1$},\\[8pt]
(a;q)_\infty & = & (1-a) (1-aq)  (1-aq^2) \cdots,\\[8pt]
 (a_1,a_2,\ldots,a_k;q)_n&=&(a_1;q)_n(a_2;q)_n\cdots (a_k;q)_n.
\end{eqnarray*}
An $_r\phi_{s}$ {\it basic hypergeometric  series} is defined by
\begin{align}\label{phi}
_r\phi_{s}\left[{a_1,a_2,\ldots,a_r \atop
b_1,b_2,\ldots,b_{s}};q,z\right]:=
\sum_{n=0}^\infty\frac{(a_1,a_2,\ldots,a_r;q)_n} {(q,b_1,\ldots,
b_{s};q)_n}\left[(-1)^nq^{n\choose 2}\right]^{1+s-r}z^n,
\end{align}
where $q\neq0$ when $r>s+1$. Further,  an $_r\psi_s$ {\it bilateral
basic hypergeometric  series} is defined by
\begin{align}\label{psi}
_r\psi_s\left[ {a_1,a_2,\ldots,a_r \atop b_1,b_2,\ldots,b_s};q,z
\right]:=\sum_{n=-\infty}^\infty\frac{(a_1,a_2,\ldots,a_r;q)_n}
{(b_1,b_2,\ldots,b_s;q)_n}\left[(-1)^nq^{n\choose
2}\right]^{s-r}z^n.
\end{align}
We assume that $q$, $z$ and the parameters are such that each
 term of the series is well-defined. We say that an $_r\phi_s$ series
terminates if one of its numerator parameters is of the form
$q^{-m}$ with $m=0,1,2,\ldots,$ and $q\neq 0$. Otherwise, we say
that the series $_r\phi_s$ is nonterminating \cite{GR2004}.

   For the ordinary  nonterminating hypergeometric identities,
Gessel \cite{Gessel} and Koornwinder \cite{Koor} provided computer
proofs of Gauss' summation formula and Saalsch\"utz' summation
formula by means of a combination of Zeilberger's algorithm and
asymptotic estimates. Vidunas
 \cite{Vidunas02} (see also Koepf \cite{Koepf03} and Koornwinder
\cite{Koorweb}) presented a method to evaluate $_2F_1\big({a,b \atop
c}\big|-1\big)$   when $c-a+b$ is an integer.
 Recently, Chen, Hou
and Mu  \cite{CHM2}  developed an approach to proving nonterminating
basic hypergeometric identities based on the $q$-Zeilberger
algorithm \cite{Koor0}. In this paper we will show how to apply the
$q$-WZ method to prove nonterminating basic hypergeometric summation
formulas by finding the  $q$-WZ pairs. We will present some examples
including
 the $q$-Gauss sum, the
$_6\phi_5$ sum, Ramanujan's $_1\psi_1$ sum and Bailey's
 $_6\psi_6$ sum  \cite{GR2004}.

\section{ The Andrews-Warnnar Identities}

In  this paper, we  give telescoping proofs of two identities on
partial theta functions of  Andrews and Warnnar \cite[Theorems 1.1
and  3.1]{GS06}:
    \begin{eqnarray}
\bigg(\sum_{n=0}^\infty(-1)^na^nq^{n\choose2
}\bigg)\bigg(\sum_{n=0}^\infty(-1)^nb^nq^{n\choose2
}\bigg)&=&(q,a,b;q)_\infty\sum_{n=0}^\infty\frac{(abq^{n-1};q)_n}
{(q,a,b;q)_n}q^n,
\label{Andrews-Warnaar-1}\\[8pt]
 1+\sum_{n=1}^\infty(-1)^nq^{n\choose
2}(a^n+b^n)&=&(a,b,q;q)_\infty \sum_{n=0}^\infty\frac{(ab/q;q)_{2n}}
{(q,a,b,ab;q)_n}q^n. \label{Andrews-Warnaar-2}
\end{eqnarray}
As will be seen, the telescoping proofs suggest that the  approach
developed by Chen, Hou and Mu \cite{CHM2} for proving nonterminating
basic hypergeometric identities can be extended so that infinite
$q$-shifted factorials are allowed in a $q$-hypergeometric term.
This idea immediately leads to an extension of the $q$-WZ method to
identities on infinite series.

 Note that the
formula (\ref{Andrews-Warnaar-2}) is a generalization of the
well-known Jacobi's triple product identity, when $b=q/a$, we get
the Jacobi's triple product identity
\begin{eqnarray}
\sum_{n=-\infty}^\infty(-1)^na^nq^{n\choose 2}=(a,q/a,q;q)_\infty,
\label{Jacobi's triple product identity}
\end{eqnarray}
 where $|q|<1$ and $a\neq 0$.

We now describe  how to prove the two identities
\eqref{Andrews-Warnaar-1} and \eqref{Andrews-Warnaar-2} by the
telescoping method. We consider \eqref{Andrews-Warnaar-1} first. Let
\[f(a)=\left(\sum_{n=0}^\infty(-1)^nq^{n\choose 2}a^n \right)
\left(\sum_{n=0}^\infty(-1)^nq^{n\choose 2}b^n \right).\]
 Note that the second factor does not contain the parameter $a$. It is
easily verified that
\begin{align}\label{R1}
f(a)=(1-a)f(aq)+aqf(aq^2).
\end{align}
So we are led  to show that the right hand side of
\eqref{Andrews-Warnaar-1} satisfies the same recurrence relation. Of
course, one should still  verify the boundary conditions. Let
\[g(a)=\sum_{n=0}^\infty D_n(a),\ \ \mbox{where}\ \ \ D_n(a)=(q,a,b;q)_\infty
\frac{(abq^{n-1};q)_nq^n} {(q,a,b;q)_n}.\] Then it is necessary to
show that
 \begin{equation} \label{fa} g(a) -(1-a)g(aq) -aqg(aq^2)=0.
\end{equation}
Here comes the key step of finding a telescoping relation for
$D_n(a)$.
 Note
that, for any $n\geq 0$, we have
\begin{align}\label{g1}
&D_n(a) -(1-a)D_n(aq)
-aqD_n(aq^2)\nonumber\\[8pt]
=&\frac{(abq^n;q)_n(a,b,q;q)_\infty q^n}{(q,a,b;q)_n} \bigg(
\frac{1-abq^{n-1}}{1-abq^{2n-1}}
-\frac{1-a}{1-aq^n}-\frac{aq(1-abq^{2n})}
{(1-aq^{n+1})(1-aq^n)(1-abq^n)} \bigg) \nonumber\\[8pt]
=&\frac{(abq^n;q)_n(a,b,q;q)_\infty q^n}{(q,a,b;q)_n} \bigg(
\frac{a(1-q^n)(1-bq^{n-1})}{(1-aq^n)(1-abq^{2n-1})}-
\frac{aq(1-abq^{2n})} {(1-aq^{n+1})(1-aq^n)(1-abq^n)}
\bigg) \nonumber\\[8pt]
=&z_{n+1}-z_n,
\end{align}
where
\[z_n=-\frac{(1-q^n)(1-bq^{n-1})(abq^n;q)_n(q,a,b;q)_\infty
aq^n}{(1-aq^n)(1-abq^{2n-1})(q,a,b;q)_n}.\] The above relation
reveals that the infinite $q$-shifted factorial $(a,b,q;q)_\infty$
can be incorporated into the telescoping relation and this step can
be automated by the $q$-Gosper algorithm. Moreover, one sees that
infinite $q$-shifted factorials can be incorporated into the
$q$-Zeilberger algorithm so that the recent approach of Chen, Hou
and Mu \cite{CHM2} can be extended. In particular, one can make the
$q$-WZ method work for infinite  basic hypergeometric  series.

 Now, let us return our attention to the proof of
 \eqref{Andrews-Warnaar-1}.
 Clearly, $z_0=0$. It is also easily seen that
  $\lim_{n\rightarrow¡¡+\infty } z_n=0$.
 Summing \eqref{g1} over all nonnegative integers, we obtain the
 recurrence relation \eqref{fa}. In order to show that $f(a)=g(a)$,
 we will use the recurrence relation of $f(a)-g(a)$ to reach this goal.

Let $H(a)=f(a)-g(a)$. From the recurrence relations for $f(a)$ and
$g(a)$, it follows that  $H(a)$ satisfies the  recurrence
 relation
\begin{align}\label{recurrence relation1}
H(a)=(1-a)H(aq)+aqH(aq^2).
\end{align}
Iterating the above relation yields that
\begin{align}{\label{Recurrence}}
H(a)=A_nH(aq^{n+1})+B_nH(aq^{n+2}),
\end{align}
 where $A_n$ and $B_n$ are given by
\[A_0=(1-a),\  B_0=aq,\   A_1=(1-a)(1-aq)+aq,\  B_{1}=(1-a)aq^2,\]
and
\[A_{n+1}=(1-aq^{n+1})A_n+aq^{n+1}A_{n-1},
\  B_{n+1}=aq^{n+2}A_n, \ \
 n\geq 1. \]
Hence we have
\[
A_{n+1}-A_{n}=-aq^{n+1}(A_n-A_{n-1}),
\]
which implies that
\begin{align*}
|A_{n+1}-A_{n}|  &=|(-1)^na^nq^{(n+1)!}||(A_1-A_0)|\\[8pt]
&\leq |a^nq^{(n+1)!}|\left(|A_1|+|A_0|\right).
\end{align*}
So, for fixed $a$ and $|q|<1$, the limit $\lim\limits_{n\rightarrow
+\infty}A_n$ exists. Since $B_{n+1}=aq^{n+2}A_n$,   the limit
$\lim\limits_{n\rightarrow +\infty}B_n$ also exists.  Again, by the
relation \eqref{Recurrence}, we find
 \[
 H(a)=H(0)\left(\lim_{n\rightarrow +\infty}A_n+\lim_{n\rightarrow
 +\infty}B_n\right).
 \]
 It remains to show that $H(0)=0$, that is,
 \begin{equation}\label{h0}
 \sum_{n=0}^\infty(-1)^nb^nq^{n\choose 2}
=(q,b;q)_\infty\sum_{n=0}^\infty\frac{q^n}{(q,b;q)_n}.
 \end{equation}
 Indeed, the above relation is a limiting case of Heine's
 transformation of $_2\phi_1$. For completeness, we give a derivation based on
 Euler's identities:
\begin{eqnarray}\label{identity-1}
(q,b;q)_\infty\sum_{n=0}^\infty\frac{q^n}{(q,b;q)_n} &=&
(q;q)_\infty\sum_{m=0}^\infty
\frac{q^m}{(q;q)_m}\sum_{n=0}^\infty\frac{q^{n\choose
2}(-bq^m)^n}{(q;q)_n}\nonumber\\[6pt]
&=&(q;q)_\infty\sum_{n=0}^\infty\frac{(-1)^nb^nq^{n\choose
2}}{(q;q)_n}\sum_{m=0}^\infty\frac{(q^{n+1})^m}{(q;q)_m}\nonumber
\\[6pt]
&=& \sum_{n=0}^\infty(-1)^nb^nq^{n\choose 2}.
\end{eqnarray}
Thus, we  have verified that $H(a)=0$.  This  completes the proof.
\qed

We remark that once the recurrence  relation \eqref{recurrence
relation1} is derived, one can also use the theorem of Chen, Hou and
Mu \cite[Theorem 3.1]{CHM2} to prove the existence of the limits of
$A_n$ and $B_n$.

We next present to a telescoping proof of \eqref{Andrews-Warnaar-2}.
Let
\[f(a)=1+\sum_{n=1}^\infty(-1)^nq^{n\choose 2}(a^n+b^n).\]
It is easily seen that
\begin{align}\label{R2}
(1+aq)f(a)-(1-a^2q)f(aq)-(aq+a^2q)f(aq^2)=(q-1)a.
\end{align}
 Let
\[g(a)=\sum_{n=0}^\infty D_n(a), \ \ \mbox{where}
\ \ \ D_n(a)=(q,a,b;q)_\infty
\frac{(ab/q;q)_{2n}q^n}{(q,a,b,ab;q)_n}.\] It will be shown that
\begin{align}\label{ga}
(1+aq)g(a)-(1-a^2q)g(aq)-(aq+a^2q)g(aq^2)=(q-1)a.
\end{align}
 Since
 \allowdisplaybreaks
\begin{align*}
&\frac{q^n-abq^{n-1}}{1-abq^{2n-1}}
-\frac{(1-a^2q)(1-ab)q^n}{(1+aq)(1-aq^n)
(1-abq^n)}\\[8pt]
&\qquad-\frac{(a^2q+aq)(1-abq^{2n})(1-abq)q^n}
{(1+aq)(1-aq^n)(1-aq^{n+1})
(1-abq^n)(1-abq^{n+1})}\\[8pt]
&\ =\frac{(1-abq^{2n})(-1+q+abq^{n+1}+a^2bq^{n+2}
-aq^{n+2}-q^{n+1}-a^2bq^{2n+2} +a^2bq^{2n+3})a}{
(1-aq^{n+1})(1-aq^n)(1-abq^{n})(1+aq)(1-abq^{n+1})}\\[8pt]
&\ -\frac{
(-1+q+abq^n+a^2bq^{n+1}-aq^{n+1}-q^n-a^2bq^{2n}+a^2bq^{2n+1})
a(1-bq^{n-1})(1-q^n)} {(1-aq^n)(1-abq^{2n-1})(1+aq)(1-abq^n)},
\end{align*}
 multiplying both sides by
$$\frac{(ab;q)_{2n}(a,b,q;q)_\infty}{(q,a,b,ab;q)_n},$$
we deduce that
\begin{align}\label{g2}
&D_n(a) -\frac{(1-a^2q)}{1+aq} D_n(aq)
-\frac{(a^2q+aq)}{1+aq}D_n(aq^2) =z_{n+1}-z_{n},
\end{align}
where
\begin{align*}
z_n=&\frac{(-1+q-aq^{n+1}-q^n+a^2bq^{n+1}+abq^n-a^2bq^{2n}
+a^2bq^{2n+1})a}{(1-aq^n)(1-abq^{2n-1})}\\[8pt]
 &\qquad\qquad\times\frac{(1-bq^{n-1})(1-q^n)(ab;q)_{2n}(q,a,b;q)_\infty}
{(1+aq)(1-abq^n)(ab;q)_n(q,a,b;q)_n}.
\end{align*}
Clearly, $z_0=0$  and $\lim_{n\rightarrow +\infty}z_{n+1}
=\frac{(q-1)a}{1+aq}$.
 Summing \eqref{g2} over all  nonnegative integers, we obtain the
 recurrence relation \eqref{ga}.

  Let
$H(a)=f(a)-g(a)$. Then $H(a)$ satisfies the following recurrence
 relation
\begin{align}\label{recurrence relation2}
H(a)=\frac{1-a^2q}{1+aq}H(aq)+ \frac{aq+a^2q}{1+aq}H(aq^2).
\end{align}
By iteration, we obtain
\begin{align}{\label{Recurrence2}}
H(a)=A_nH(aq^{n+1})+B_nH(aq^{n+2}),
\end{align}
 where $A_n$ and $B_n$ are given by
\begin{align*}
&A_0=\frac{1-a^2q}{1+aq},\ \ \ A_1=\frac{(1-a^2q^3)(1-a^2q)}
{(1+aq)(1+aq^2)}+\frac{aq+a^2q}{1+aq},\\[8pt]
&B_0=\frac{aq+a^2q}{1+aq},\ \ \
B_1=\frac{(1-a^2q)(aq^2+a^2q^3)}{(1+aq)(1+aq^2)},
\end{align*}
and  for $n \geq 1$,
\begin{eqnarray}
 A_{n+1}& = & \frac{1-a^2q^{2n+3}}{1+aq^{n+2}}A_{n}
+\frac{aq^{n+1}+a^2q^{2n+1}}{1+aq^{n+1}}A_{n-1},\label{Formula-1} \\[6pt]
B_{n+1}& =& \frac{aq^{n+2}+a^2q^{2n+3}}{1+aq^{n+2}}A_n.
\label{Formula-2}
\end{eqnarray}
It follows from \eqref{Formula-1} that
\begin{align}{\label{relation1}}
A_{n+1}-\frac{-aq^{n+2}-a^2q^{2n+3}}{1+aq^{n+2}}A_n=
A_n-\frac{-aq^{n+1}-a^2q^{2n+1}}{1+aq^{n+1}}A_{n-1}. \ \ \ \ n\geq1.
\end{align}
For given  $a$ and $|q|<1$, we have
\[\lim_{n\rightarrow +\infty}\frac{-aq^{n+2}-a^2q^{2n+3}}
{1+aq^{n+2}}=0 .\] Now, for $\varepsilon_0=\frac{1}{2}$, there
exists an integer $N_0>0$, depending on  $a$ and $q$,
 such that for any integer $n\geq N_0$, we have
\begin{align}{\label{relation2}}
\left|\frac{-aq^{n+2}-a^2q^{2n+3}}{1+aq^{n+2}}
\right|<\varepsilon_0=\frac{1}{2}.
\end{align}
Therefore,  given  $a$ and $|q|<1$, for any integer $n\geq N_0$, we
have \allowdisplaybreaks
\begin{align*}
&|A_{n+1}-A_{N_0}|\\[8pt]
=&{\bigg|} A_{n+1} -\frac{-aq^{n+2}-a^2q^{2n+3}}{1+aq^{n+2}}A_n+
\frac{-aq^{n+2}-a^2q^{2n+3}}{1+aq^{n+2}}
(A_n-\frac{-aq^{n+1}-a^2q^{2n+1}}{1+aq^{n+1}}A_{n-1})\\[8pt]
+&\cdots \\[8pt]
+&\frac{(-aq^{n+2}-a^2q^{2n+3}) \cdots(-aq^{N_0+3}-a^2q^{2N_0+5})}
{(1+aq^{n+2})\cdots(1+aq^{N_0+3})}(A_{N_0+1}-
\frac{-aq^{N_0+2}-a^2q^{2N_0+3}}{1+aq^{N_0+2}}A_{N_0})\\[8pt]
+&\frac{(-aq^{n+2}-a^2q^{2n+3})
\cdots(-aq^{N_0+3}-a^2q^{2N_0+5})(-aq^{N_0+2}-a^2q^{2N_0+3})}
{(1+aq^{n+2})\cdots(1+aq^{N_0+3})(1+aq^{N_0+2})}A_{N_0}-A_{N_0}
 {\bigg|}\\[8pt]
 \leq & \left|A_{n+1} -\frac{-aq^{n+2}-a^2q^{2n+3}}
 {1+aq^{n+2}}A_n\right|
 +\left|\frac{-aq^{n+2}-a^2q^{2n+3}}{1+aq^{n+2}}\right|
\left|(A_n-\frac{-aq^{n+1}-a^2q^{2n+1}}{1+aq^{n+1}}
A_{n-1})\right|\\[8pt]
+&\cdots \\[8pt]
+&\left|\frac{(-aq^{n+2}-a^2q^{2n+3})
\cdots(-aq^{N_0+3}-a^2q^{2N_0+5})}
{(1+aq^{n+2})\cdots(1+aq^{N_0+3})}\right|\left|(A_{N_0+1}-
\frac{-aq^{N_0+2}-a^2q^{2N_0+3}}
{1+aq^{N_0+2}}A_{N_0})\right|\\[8pt]
+&\left|\frac{(-aq^{n+2}-a^2q^{2n+3})
\cdots(-aq^{N_0+3}-a^2q^{2N_0+5})(-aq^{N_0+2}-a^2q^{2N_0+3})}
{(1+aq^{n+2})\cdots(1+aq^{N_0+3})(1+aq^{N_0+2})}A_{N_0}\right|
+|A_{N_0}|\\[8pt]
\leq &(1+\frac{1}{2}+\frac{1}{2^2}+\cdots +\frac{1}{2^{n-N_0}})
\left|A_{N_0+1}-
\frac{-aq^{N_0+2}-a^2q^{2N_0+3}}{1+aq^{N_0+2}}A_{N_0}\right|
+(1+\frac{1}{2^{n+1-N_0}})|A_{N_0}|\\[8pt]
\leq & 3|A_{N_0+1}|+3\left|\frac{-aq^{N_0+2}-a^2q^{2N_0+3}}
{1+aq^{N_0+2}}A_{N_0}
\right|+\frac{3}{2}|A_{N_0}|\\[8pt]
 \leq & 3|A_{N_0+1}|+3|A_{N_0}|.
\end{align*}
Here we have used the relations \eqref{relation1} and
\eqref{relation2}. From the above inequality, we find
\[
|A_{n+1}|\leq 3|A_{N_0+1}|+4|A_{N_0}|, \ \ \ \ \ \ n\geq N_0-1.
\]
Thus, for  $n\geq N_0+1$, we have
\begin{align*}
|A_{n+1}-A_n|&=\left|\frac{-aq^{n+2}-a^2q^{2n+3}}{1+aq^{n+2}}A_n+
\frac{aq^{n+1}+a^2q^{2n+1}}{1+aq^{n+1}}A_{n-1}\right|\\[8pt]
&\leq \left(\left|\frac{-aq^{n+2}-a^2q^{2n+3}}{1+aq^{n+2}}\right| +
\left|\frac{-aq^{n+1}-a^2q^{2n+1}}{1+aq^{n+1}}\right|\right)
(3|A_{N_0+1}|+4|A_{N_0}|).
\end{align*}
It follows that the limit $\lim\limits_{n\rightarrow +\infty}A_n$
exists. Since
\[ B_{n+1}=\frac{aq^{n+2}+a^2q^{2n+3}}{1+aq^{n+2}}A_n,\]
 the limit $\lim\limits_{n\rightarrow +\infty}B_n$ also exists. By the
relation \eqref{Recurrence2}, we deduce that
 \[
 H(a)=H(0)(\lim_{n\rightarrow +\infty}A_n+\lim_{n\rightarrow
 +\infty}B_n).
 \]
The identity \eqref{identity-1} implies that  $f(0)=g(0)$. So we
have $H(a)=0$. This completes the proof. \qed

We also note that once the recurrence  relation \eqref{recurrence
relation2} is derived, one may assume that  $|a|<1$ and may use the
the  theorem in Chen, Hou and Mu \cite[Theorem 3.1]{CHM2}) to prove
the existence of the limits of $A_n$ and $B_n$. Moreover, we may
drop the assumption $|a|<1$ by analytic continuation.

\section{The $q$-WZ Pairs for Infinite Series}

Our approach to the $q$-WZ method for infinite series can  be
described as follows. The key step is to construct $q$-WZ pairs for
infinite sums.  Suppose that we aim to prove an identity of the
form:
\begin{align}\label{equation1}
\sum_{k=N_0}^\infty F_k(a_1,a_2,\ldots, a_t)=R(a_1,a_2,\ldots,a_t),
\end{align}
where $t$ is a positive integer, and the sum is either a unilateral
or bilateral basic hypergeometric series, namely, $N_0=0$ or
$N_0=-\infty$,  $R(a_1,a_2,\ldots,a_t)$ is either zero or a quotient
of two products
 of infinite $q$-shifted  factorials.

First, we set some parameters, say, $a_1,\ldots,a_p$, ($1\leq p\leq
t$) to $a_1q^n,\ldots, a_pq^n$, so that we get
\begin{align}\label{ Set parameters }
\sum_{k=N_0}^\infty F_k(a_1q^n,\ldots,
a_pq^n,a_{p+1},\ldots,a_t)=R(a_1q^n,\ldots,
a_pq^n,a_{p+1},\ldots,a_t).
\end{align}
If $R(a_1q^n,\ldots, a_pq^n,a_{p+1},\ldots,a_t)\neq 0$, Then set
\[
F(n,k)=\frac{F_k(a_1q^n,\ldots,
a_pq^n,a_{p+1},\ldots,a_t)}{R(a_1q^n,\ldots,
a_pq^n,a_{p+1},\ldots,a_t)}.
\]
Otherwise, we simply set
\[
F(n,k)=F_k(a_1q^n,\ldots, a_pq^n,a_{p+1},\ldots,a_t).
\]
Our goal is to show that
\begin{align}\label{Constant}
\sum_{k=N_0}^\infty F(n,k)= {\rm constant}, \ \ \ \ \ \
n=0,1,2,\ldots.
\end{align}
The constant can be determined by setting $n=0$ and setting
$a_1,a_2,\ldots, a_t$ to special values. We claim that the above
goal can be achieved by adopting the $q$-WZ method for infinite
sums.

Let us recall the boundary and limit conditions for the $q$-WZ
method. Let $f(n)$ denote the left hand side of \eqref{Constant},
i.e.,
\[ f(n)=\sum\limits_{k=N_0}^\infty F(n,k)\]
 and we aim to show that
\[ f(n)={\rm constant}\]
 for every nonnegative integer $n$. To this end, it suffices to
show that $f(n+1)-f(n)=0$ for every nonnegative integer $n$. This
can be done by finding  $G(n,k)$ such that
\begin{align}\label{WZPair}
F(n+1,k)-F(n,k)=G(n,k+1)-G(n,k).
\end{align}
A pair of functions $(F(n,k),G(n,k))$ that satisfy \eqref{WZPair} is
called a {\it $q$-WZ pair}. Once a $q$-WZ pair is found, one can
check the boundary and limit conditions to ensure that $f(n)$ equals
the claimed constant. Here are the conditions:
\begin{itemize}
\item[(C1)] For each integer $n\geq 0$,
$\lim\limits_{k\rightarrow\pm \infty}G(n,k)=0$.

\item[(C2)] For each integer $k$, the limit
\begin{align}\label{limit}
f_k=\lim_{n\rightarrow \infty}F(n,k)
\end{align}
\ \ \ \ \ \ \ \ \ \ \  \ exists and is finite.

\item[(C3)] We have $\lim\limits_{L\rightarrow
\infty}\sum\limits_{n\geq0} G(n,-L)=0$.
\end{itemize}

The WZ method can be formally stated as follows.

\begin{thm}[Wilf and Zeilberger \cite{WZ2}]
 Assume that $(F(n,k),G(n,k))$ is a WZ pair \eqref{WZPair}.
 If {\rm (C1)} holds, then we have
\begin{align}\label{Con1}
\sum_{k}F(n,k)={\rm constant},\ \ \ \ \ \ \ n=0,1,2,\ldots.
\end{align}
If {\rm(C2)} and {\rm(C3)} hold, then we have the companion identity
\begin{align}\label{Identity}
\sum_{n=0}^\infty G(n,k)=\sum_{j\leq k-1}(f_j-F(0,j)),
\end{align}
 where $f_j$ is defined by \eqref{limit}.
\end{thm}

We now explain how to compute the desired $q$-WZ pair for the
identity \eqref{equation1}.
 In fact, it can be produced by applying the
 $q$-Gasper algorithm to
$F(n+1,k)-F(n,k)$. It should be noted that
  $F(n+1,k)-F(n,k)$ is  a  $q$-hypergeometric term with respect to
   $q^k$,  even if $F(n,k)$ contains
  infinite $q$-shifted factorials such as $(aq^n;q)_\infty.$
 Obviously,  $F(n+1,k)-F(n,k)$ is a  $q$-hypergeometric term when $R(a_1,\ldots,a_t)= 0$.
 Assume that $R(a_1,\ldots,a_t)\neq0$.  Let
 \allowdisplaybreaks
 \begin{align*}
 M_1&=\frac{R(a_1q^{n+1},\ldots,
a_pq^{n+1},a_{p+1},\ldots,a_l)}{R(a_1q^n,\ldots,
a_pq^n,a_{p+1},\ldots,a_t)},\\[8pt]
M_2 &=\frac{F_{k+1}(a_1q^{n+1},\ldots,
a_pq^{n+1},a_{p+1},\ldots,a_t)}{F_k(a_1q^{n+1},\ldots,
a_pq^{n+1},a_{p+1},\ldots,a_t)},\\[8pt]
M_3&=\frac{F_{k+1}(a_1q^n,\ldots,
a_pq^n,a_{p+1},\ldots,a_t)}{F_k(a_1q^{n+1},\ldots,
a_pq^{n+1},a_{p+1},\ldots,a_t)},\\[8pt]
M_4&=\frac{F_k(a_1q^n,\ldots,
a_pq^n,a_{p+1},\ldots,a_t)}{F_k(a_1q^{n+1},\ldots,
a_pq^{n+1},a_{p+1},\ldots,a_t)}.
 \end{align*}
 Since $M_1$ is a rational function in $q^n$ and is independent of
 $k$,
 $M_2, M_3, M_4$ are all rational functions in $q^k$. Observe that
\begin{equation}\label{fm}
 \frac{F(n+1,k+1)-F(n,k+1)}{F(n+1,k)-F(n,k)}=
\frac{M_2-M_1M_3}{1-M_1M_4}
\end{equation}
is a rational function in $q^k$,  i.e.,  $F(n+1,k)-F(n,k)$ is a
$q$-hypergeometric term with respect to $q^k$. It is necessary to
mention that even if $F(n,k)$ contains infinite $q$-shifted
factorials of the form $(aq^n;q)_\infty$, the quotient \eqref{fm}
 no longer contains the $q$-shifted
factorial $(aq^n;q)_\infty$ and it is still a rational function in
$q^k$. Consequently, we can employ the $q$-Gosper algorithm to
determine whether $G(n,k)$ exists.
 Nevertheless, it is also necessary to note that $G(n,k)$
contains infinite $q$-shifted factorials if $F(n,k)$ does.

There is another way to look at the above procedure.  Suppose that
$F(n,k)$ contains an infinite $q$-shifted factorial $(a;q)_\infty$,
where $a$ is a chosen parameter for the substitution $a\rightarrow
aq^n$. If we set $G'(n,k)=R(aq^n) G(n,k)$. Then the equation
\eqref{WZPair} becomes
\[ F(n+1, k)R(aq^{n})  - F(n,k) R(aq^n) = G'(n,k+1) - G'(n,k).\]
It is evident that the infinite $q$-shifted factorial
$(aq^n;q)_\infty$ will disappear in the above equation, and one can
use the $q$-Gosper algorithm to find a $q$-WZ pair if it exists.

We now take the $q$-binomial theorem  \cite[P. 354]{GR2004} as an
example to explain the above steps:
\begin{align}\label{q-binomial theorem}
\sum_{k=0}^\infty\frac{(a;q)_k}{(q;q)_k}z^k=
\frac{(az;q)_\infty}{(z;q)_\infty},\ \ \ \ \ |z|<1.
\end{align}
In this case,  we have \[ F_k(a)= \frac{(a;q)_k}{(q;q)_k}z^k, \quad
R(a)=\frac{(az;q)_\infty}{(z;q)_\infty}. \] We choose the parameter
$a$, and substitute $a$ with $aq^n$. Then we set
\[ F(n,k)={F_k(aq^n) \over R(aq^n)} ={ (aq^n;q)_k (z;q)_\infty  \over (q;q)_k
(azq^n;q)_\infty}z^k. \] In order to find $G(n,k)$ such that
 \eqref{WZPair} holds, it is easily checked that $F(n+1, k)-F(n,k)$ is a
$q$-hypergeometric term. By examining the $q$-Gosper algorithm, one
sees that it is capable to deal with the input $F(n+1,k)-F(n,k)$, or
we can set \[ G'(n,k)= R(aq^n) G(n,k)\] and find a solution of the
equation
\begin{equation}\label{New-WZ-Equation}
(1-azq^n)\frac{(aq^{n+1};q)_k}{(q;q)_k}z^k-\frac{(aq^n;q)_k}{(q;q)_k}z^k
= G'(n,k+1)-G'(n,k).
\end{equation}
Finally, we obtain the  $q$-WZ pair
\begin{align*}
F(n,k)&=\frac{(aq^n;q)_k(z;q)_\infty}
{(q;q)_k(azq^n;q)_\infty}z^k,\\[8pt]
G(n,k)&=-\frac{(aq^n;q)_k(z;q)_\infty(a-aq^k)}
{(q;q)_k(azq^n;q)_\infty(1-aq^n)}q^nz^k.
\end{align*}
If $|z|<1$,  it is easy to see that $F(n,k)$ and $G(n,k)$ satisfy
the conditions (C1), (C2) and (C3). By \eqref{Con1},
\[
\sum_{k=-\infty}^\infty F(n,k)=\sum_{k=0}^\infty F(n,k)={\rm
constant}, \ \ \ \ \ n=0,1,2,\ldots.
\]
Setting $z=0$ yields that the constant equals 1. Setting $n=0$,  we
have
\[
\sum_{k=0}^\infty F(0,k)={\rm constant}=1.
\]
 By
\eqref{Identity}, we get the companion identity of \eqref{q-binomial
theorem}
\begin{align*}
\sum_{j=0}^k\frac{(a;q)_j}{(q;q)_j}z^j=
(az;q)_\infty\sum_{j=0}^k\frac{z^j}{(q;q)_j}+
\frac{az^{k+1}(a;q)_{k+1}}{(q;q)_{k}}\sum_{n=0}^\infty
\frac{(az;q)_n(aq^{k+1};q)_n}{(a;q)_{n+1}}q^n.
\end{align*}

We now give a few more examples.

 \noindent {\bf
Example 3.1.} The $q$-Gauss sum \cite[P. 354]{GR2004}:
\begin{align}\label{Gauss}
\sum_{k=0}^\infty
\frac{(a,b;q)_k}{(q,c;q)_k}\left(\frac{c}{ab}\right)^k
=\frac{(c/a,c/b;q)_\infty} {(c,c/ab;q)_\infty},\ \ \ \ |c/ab|<1.
\end{align}

By computation, we obtain the following $q$-WZ pair
\begin{align*}
F(n,k)&=\frac{(b,aq^n;q)_k(c/ab,cq^n;q)_\infty}
{(q,cq^n;q)_k(c/a,cq^n/b;q)_\infty}
\left(\frac{c}{ab}\right)^k,\\[8pt]
G(n,k)&=-\frac{(a-aq^k)(b,aq^n;q)_k(c/ab,cq^n;q)_\infty}
{(1-aq^n)(q,cq^n;q)_k(c/a, cq^n/b;q)_\infty}
\left(\frac{c}{ab}\right)^kq^n.
\end{align*}
If $|c/ab|<1$, it is easy to verify that $F(n,k)$ and $G(n,k)$
satisfy the conditions (C1), (C2) and
 (C3).  By \eqref{Con1}, we have
\[
\sum_{k=-\infty}^\infty F(n,k)=\sum_{k=0}^\infty F(n,k)= {\rm
constant}, \ \ \ \ n=0,1,2,\ldots.
\]
Setting $c=0$ and $n=0$, we find that the constant equals 1, and  we
have
\[
\sum_{k=0}^\infty F(0,k)={\rm constant}=1.
\]
After simplification, we obtain the identity \eqref{Gauss}.

 By
\eqref{Identity}, we obtain the companion identity of \eqref{Gauss}
\begin{align}\label{Gauss-c}
-\sum_{n=0}^\infty \frac{(a-aq^k)(b,aq^n;q)_k(c/ab,cq^n;q)_\infty}
{(1-aq^n)(q,cq^n;q)_k(c/a, cq^n/b;q)_\infty}
\left(\frac{c}{ab}\right)^kq^n =\sum_{j\leq k-1}(f_j-F(0,j)),
\end{align}
where
\[
f_j=\lim_{n\rightarrow \infty}F(n,j)=\frac{(b;q)_j(c/ab;q)_\infty}
{(q;q)_j(c/a;q)_\infty}\left(\frac{c}{ab}\right)^j,
\]
which can be restated as
\begin{align*}
\sum_{j=0}^k
\frac{(a,b;q)_j}{(q,c;q)_j}\left(\frac{c}{ab}\right)^j=&
\frac{(c/b;q)_\infty}{(c;q)_\infty}
\sum_{j=0}^k\frac{(b;q)_j}{(q;q)_j}
\left(\frac{c}{ab}\right)^j\\[8pt]
&+\frac{(a,b;q)_{k+1}c^{k+1}}{(q;q)_k(c;q)_{k+1}a^kb^{k+1}}
\sum_{n=0}^\infty\frac{(aq^{k+1},c/b;q)_n}
{(a;q)_{n+1}(cq^{k+1};q)_n}q^n.
\end{align*}

 \noindent{\bf {Example 3.2.}}
 The  $_6\phi_5$
 summation formula \cite[P. 356]{GR2004}:
\begin{align}\label{65}
\sum_{k=0}^\infty &
\frac{(1-aq^{2k})(a,b,c,d;q)_k}{(1-a)(q,aq/b,aq/c,aq/d;q)_k}
\left(\frac{aq}{bcd}\right)^k \nonumber\\[8pt]
&\qquad\qquad =\frac{(aq,aq/bc,aq/bd,aq/cd;q)_\infty}
{(aq/b,aq/c,aq/d,aq/bcd;q)_\infty},\ \ \ \ \ \ \ |aq/bcd|<1.
\end{align}

We get the following $q$-WZ pair:
\begin{align*}
F(n,k)=&\frac{(1-aq^{n+2k})(c,d, aq^n,bq^n;q)_k}
{(1-aq^n)(q,aq/b,aq^{n+1}/c,aq^{n+1}/d;q)_k}\\[8pt]
&\qquad \times\frac{(aq/b,aq/bcd,aq^{n+1}/c,aq^{n+1}/d;q)_\infty}
{(aq/bc,aq/bd,aq^{n+1},aq^{n+1}/cd;q)_\infty}
\left(\frac{aq}{bcd}\right)^k,\\[8pt]
G(n,k)=&\frac{(c,d;q)_k(a/b,a/bcd;q)_\infty}
{(q,a/b;q)_k(aq^n,aq^n/cd;q)_\infty}\\[8pt]
&\qquad\times \frac{(aq^n,bq^n;q)_k(aq^n/c,aq^n/d;q)_\infty}
{(aq^n/c,aq^n/d;q)_k(a/bd,a/bc;q)_\infty
}\\[8pt]
&\qquad\times \frac{(a-bc)(a-bd)(aq^n-cd)(1-q^k)}{(a-bcd)(bq^n-1)
(aq^{n+k}-c)(aq^{n+k}-d)} \left(\frac{aq}{bcd}\right)^kq^n.
\end{align*}
It is easily seen that $F(n,k)$ and $G(n,k)$ satisfy the conditions
(C1), (C2) and (C3). Therefore, by \eqref{Con1}, we have
$\sum\limits_{k=0}^\infty F(n,k)$ is a constant. Setting $n=0$ and
$a=0$, we find that the constant equals 1. Thus we have
\[ \sum_{k=0}^\infty F(0,k)={\rm constant}=1,\]
 which is nothing but \eqref{65}. Since
\[
f_k=\frac{(c,d;q)_k}{(q,aq/b;q)_k}\frac{(aq/b,aq/bcd;q)_\infty}
{(aq/bc,aq/bd;q)_\infty}\left(\frac{aq}{bcd}\right)^k
\]
and
\[
F(0,j)=\frac{(1-aq^{2j})(a,b,c,d;q)_j
(aq/b,aq/c,aq/d,aq/bcd;q)_\infty}
{(1-a)(q,aq/b,aq/c,aq/d;q)_j(aq,aq/bc,aq/bd,aq/cd;q)_\infty}
\left(\frac{aq}{bcd}\right)^j,
\]
 by \eqref{Identity},  we obtain the
companion identity
\begin{align*}
\sum_{j=0}^k&\frac{(1-aq^{2j})(a,b,c,d;q)_j }
{(1-a)(q,aq/b,aq/c,aq/d;q)_j}
\left(\frac{aq}{bcd}\right)^k\\[8pt]
&\qquad\qquad=\frac{(aq,aq/cd;q)_\infty}{(aq/c,aq/d;q)_\infty}
\sum_{j=0}^k\frac{(c,d;q)_j}
{(q,aq/b;q)_j}\left(\frac{aq}{bcd}\right)^j\\[8pt]
&\qquad\qquad+\frac{b(aq;q)_k(b,c,d;q)_{k+1}}
{(q,aq/b;q)_k(aq/c,aq/d;q)_{k+1}}
\left(\frac{aq}{bcd}\right)^{k+1}\\[8pt]
&\qquad\qquad\times\sum_{n=0}^\infty
\frac{(aq/cd;q)_n(aq^{k+1},bq^{k+1};q)_n}{(b;q)_{n+1}
(aq^{k+2}/c,aq^{k+2}/d;q)_{n}}q^n.
\end{align*}

\noindent{\bf Example 3.3.}  Ramanujan's $_1\psi_1$ sum \cite[P.
357]{GR2004}
\begin{align}\label{11}
_1\psi_1(a;b;q,z)=
\frac{(q,b/a,az,q/az;q)_\infty}{(b,q/a,z,b/az;q)_\infty}, \ \ \
|b/a|<|z|<1.
\end{align}

In this case, we find that
\begin{align*}
F(n,k)&=\frac{(aq^n;q)_k(z,b/az,bq^n,q^{1-n}/a;q)_\infty}
{(bq^n;q)_k(q,b/a,azq^n,q^{1-n}/az;q)_\infty}z^k,\\[8pt]
G(n,k)&=\frac{(z,b/az,bq^n,q^{-n}/a;q)_\infty(aq^n;q)_k(1-azq^n)}
{(q,b/a,azq^n,q^{-n}/az;q)_\infty(bq^n;q)_k(z-azq^n)}z^k.
\end{align*}
If $|b/a|<|z|<1$, utilizing the following identity
\begin{align}{\label{Trans}}
(a;q)_{-n}=\frac{(-q/a)^nq^{n\choose 2}}{(q/a;q)_n}, \ \ \ \
n=0,1,2,\ldots,
\end{align}
we can verify that $G(n,k)$ satisfies the condition (C1). It follows
that
\begin{align} \label{Ra1}
\sum_{k=-\infty}^\infty F(n,k)= {\rm constant}, \ \ \ \ \
n=0,1,2,\ldots.
\end{align}
Setting $n=0$,  $b=q$ and utilizing  the $q$-binomial theorem
\eqref{q-binomial theorem}, we see that the constant equals 1.
Setting $n=0$, we obtain the identity \eqref{11}. However, we note
  that the conditions for the companion identity do not hold in this case.

\noindent{\bf Example 3.4.} Bailey's $_6\psi_6$ summation formula
\cite[P. 357]{GR2004}:
\begin{align}\label{66}
\sum_{k=-\infty}^\infty&\frac{(1-aq^{2k})(b,c,d,e;q)_k}
{(1-a)(aq/b,aq/c,aq/d,aq/e;q)_k}
\left(\frac{a^2q}{bcde}\right)^k \nonumber\\[8pt]
&\qquad\qquad
=\frac{(aq,aq/bc,aq/bd,aq/be,aq/cd,aq/ce,aq/de,q,q/a;q)_\infty}
{(aq/b,aq/c,aq/d,aq/e,q/b,q/c,q/d,q/e,a^2q/bcde;q)_\infty}.
\end{align}

We obtain the following $q$-WZ pair:
\begin{align*}
F(n,k)=&\frac{(1-aq^{n+2k})(d,e,bq^n,cq^n;q)_k (aq/b,aq/c;q)_\infty}
{(1-aq^n)(aq/b,aq/c,aq^{n+1}/d,aq^{n+1}/e;q)_k
(aq/bd,aq/be;q)_\infty}\\[8pt]
& \qquad\times \frac{(q/d,q/e,a^2q/bcde,aq^{n+1}/d,
aq^{n+1}/e,q^{1-n}/b,q^{1-n}/c;q )_\infty}
{(q,aq/cd,aq/ce,aq^{n+1},aq^{n+1}/de,
q^{1-n}/a,aq^{1-n}/bc;q)_\infty} \left(\frac{a^2q}
{bcde}\right)^k,\\[8pt]
G(n,k)=&\frac{(d,e,bq^n,cq^n;q)_k(a/b,a/c,1/e,a^2/bcde,
1/d;q)_\infty}{(a/b,a/c,aq^n/d,aq^n/e;q)_k
(q,a/bd,a/be,a/cd,a/ce;q)_\infty }\\[8pt]
&\qquad\times \frac{(aq^n/d,aq^n/e,q^{-n}/b ,q^{-n}/c;q)_\infty
(-1+aq^n)} {
(aq^n,aq^n/de,aq^{-n}/bc,q^{-n}/a;q)_\infty(1-bq^n) (1-cq^n)}\\[8pt]
&\qquad\times\frac{(a-bd)(a-be)(a-cd)(a-ce)(aq^n-de)q^n}
{(aq^{n+k}-d)(aq^{n+k}-e)(a-ad)(1-e)(a^2-bcde)} \left(\frac{a^2q}
{bcde}\right)^k.
\end{align*}
Since $|a^2q/bcde|<1$, from  the identity \eqref{Trans} it follows
that $G(n,k)$ satisfies the condition (C1). By \eqref{Con1}, we find
\begin{align} \label{Ra2}
\sum_{k=-\infty}^\infty F(n,k)= {\rm constant}, \ \ \ \ \
n=0,1,2,\ldots.
\end{align}
In order to determine the constant, we set $n=0$ and $b=a$. From the
 $_6\phi_5$  summation formula  \eqref{65}, we see that
 the constant equals
\begin{eqnarray*}
\sum_{k=-\infty}^\infty F(0,k) & =
&\sum_{k=0}^\infty\frac{(1-aq^{2k})(a,c,d,e;q)_k}
{(1-a)(aq/c,aq/d,aq/e;q)_k
}\\[8pt]
& & \quad \times \frac{(aq,aq/cd,aq/ce,aq/de;q)_\infty}
{(aq/c,aq/d,aq/e,aq/cde;q)_\infty}\left(\frac{aq}{cde}\right)^k=1,
\end{eqnarray*}
 which can be restated as \eqref{66}. Nevertheless, we note that
the conditions for the companion identity do not hold in this case.

\vspace{1cm}
 \noindent{\bf Acknowledgments.} The authors would like to thank George E. Andrews, Q.H.
  Hou and Doron Zeilberger for valuable comments.  This work was supported by  the 973
Project, the PCSIRT Project of the Ministry of Education, the
Ministry of Science and Technology, and the National Science
Foundation of China.

\end{document}